I wold greatly appreciate some help with this word problem that I have been working on.
Here is the question: Rose's garden is in the shape of a trapezoid. If the height of the trapezoid is 16m, one base is 20m, and the area is 224m^, find the length of the other base.
What I have for the equation so far is this; L=1/2h(B+A) does this make any sense on what I have so far

Answer by checkley71(8403)

[(20+B)/2]*16=224
(20+B)*8=224
160+8B=224
8B=224-160
8B=64
B=64/8
B=8 ANSWER FOR THE SECOND BASE MEASUREMENT.
PROOF
(20+8)/2*16=224
(28/2)16=224
28*8=224
224=224